The lengths (in minutes) of a random selection of popular children’s animated films are listed below. Estimate the true mean length of all children’s animated films with 95% confidence.


91 85 77 94 80 90 78 100 78 76



a. Sample mean:


b. Sample standard deviation:


c. Margin of error:


d. Confidence interval: __________ < mu< __________

1
Expert's answer
2021-11-04T15:54:42-0400

a. Sample mean:

"\mu = \frac{91+85+...+78+76}{10} = 84.9"

b. Sample standard deviation:

"s = \sqrt{\frac{(91-84.9)^2 +(85-84.9)^2 +….+(78-84.9)^2 +(76-84.9)^2}{10-1}} = 8.399"

c. Margin of error "= Z_c \times s"

"ME = 1.96 \times 8.399 = 16.46"

d. Confidence interval: 84.9 – 16.46 < mu< 84.9 +16.46

68.44 < mu < 101.36